OpenStudy (anonymous):

Divide (10 × 10^17) ÷ (100 × 10^13). Express answer in scientific notation

OpenStudy (anonymous):

\[\frac{10*10^{17}}{100*10^{13}}=\frac{1*10^{18}}{1*10^{15}}=1*10^3\]